OpenStudy (lilai3):

PLEASE HELP!!! WILL REWARD MEDAL/ FAN/ WRITE TESTIMONY, ETC. ETC.!!! Define a variable, write an inequality, and solve each problem. Then interpret your solution. There are 3,500 species of bees and more than 600,000 species of insects. How many species of insects are not bees? Please help me and explain. Thank you.~

OpenStudy (anonymous):

bee = 3500 other insects + bees > 600,000 other insects + 3500 > 600,00 other insects > 596,500

OpenStudy (shamil98):

just a wild guess x = 3500 y > 600000 let z be the insects that are not bees \[\huge z \le 596500\] i think.

OpenStudy (shamil98):

wrong sign* \[\huge z \ge 596500\]

OpenStudy (ranga):

just "greater than" and not "greater than or equal" since "more than 600,000"

OpenStudy (shamil98):

oh

OpenStudy (shamil98):

\[\huge z > 596500\] then

OpenStudy (lilai3):

wait..what? i'm kinda confused...

OpenStudy (ranga):

non-bees = x bees = 3500 x + 3500 > 600,000 x > 596,500

OpenStudy (lilai3):

wait...i got n+3,500<600,000...

OpenStudy (ranga):

Insects = non-bees + bees = n + 3,500 "there are more than 600,000 species of insects" Insects > 600,000 n + 3,500 > 600,000

OpenStudy (lilai3):

Wait... I thought the sign was the other way...<....because it stated that there was "more than 600,000", not n+3,500 was more....wha-?...

OpenStudy (lilai3):

hello, @ranga @shamil98 ? hello? help, please?

OpenStudy (ranga):

There are MORE THAN 600,000 species of insects. For this problem, the insects are divided into two groups: bees and non-bees. The non-bees includes all insects other than bees. So insects = bees + non-bees > 600,000 ---- (1) Assume there are n species of non-bees. There are 3500 species of bees. Plug those values into (1): So insects = 3,500 + n > 600,000

OpenStudy (lilai3):

oh. oh. thank you for clarifying. my mistake, i'm sorry. thanks for teaching me, @ranga ! hope you enjoyed the medal i gave you! ;D

OpenStudy (ranga):

You are welcome.